OpenStudy (obeymesenpai):

Jeffrey is training to run a marathon. On Monday he ran and walked a total of 11 miles. If he ran 2 miles more than twice as far as he walked, how many miles did he run? A) 3 Miles B) 5 Miles C) 8 Miles D)10 Miles

OpenStudy (tgstudios):

well, start of with 11 divided by 3

OpenStudy (tgstudios):

hold on I have to go, somebody else help!

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)

OpenStudy (legendofmath):

I GOTCHU FAM

OpenStudy (legendofmath):

So first you need to find out how much he walked and ran. You could use an equation like w+r=11

OpenStudy (legendofmath):

then, it says he ran 2 miles more than twice as far as he walked. so you would find how much he ran: Ran= 2r+2 After you have this you can plug it into the equation before (w+r=11): w+(2w+2)= 11 and frm here you can do the math and get how much he walked. w+2w+2 3w+2=11 -2 -2 3w=9 -- -- 3 3 w=3

OpenStudy (legendofmath):

So now you know how much he walked. And you can plug this into the original equation (w+r=11) and figure out how much he walked: 3+r=11 -3 -3 r=8 so he ran 8 miles (dayum thats alot) Good luck with the other problems :D

OpenStudy (triciaal):

@legendofmath just a note: a marathon is 26 miles (I am not a runner either)